OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = −3x + 12 and passes through the point (−1, 6).

OpenStudy (anonymous):

y = one thirdx + 7 y = −3x + 3 y = one thirdx + 3 y = −3x + 7 these are the choices

OpenStudy (jdoe0001):

same as the one before, just that, parallel lines have EXACTLY the same slope, so the slope of => y = −3x + 12 is -3 so any parallel line to it will also have a slope of -3 so => \(\bf \large y-y_1=m(x-x_1)\) m = slope (-1, 6) = \(\bf \large (x_1, y_1)\)